# MAVERIX: Multimodal Audio-Visual Evaluation Reasoning IndeX

### v1

@@ D.6 Common Error Modes
To better understand the limitations of the evaluated models, we analyze common error patterns observed across the benchmark. We find that model failures often arise from modality-specific weaknesses, such as misinterpretation of complex visual cues, inability to process nuanced audio signals, or challenges in aligning multimodal information with nuanced context. Common failure modes are showcased in Fig. D.2,  D.3,  D.4,
@@
Figure D.2: Error analysis showing that o1 fails to correctly answer the question when the audio cannot be transcribed into text-based subtitles, leading to an incorrect connection between the cityscape video and the instrumental music.
@@
Figure D.3: Error analysis showing inconsistent responses from o1, where it failed to accurately identify Hika-chan, when the question was rephrased slightly.
@@
Figure D.4: Error analysis highlighting o1’s failure in correctly synchronizing the audio with visual cues, leading to misinterpretation of the woman picking up tree ornaments.
@@
Figure D.5: Error analysis highlighting o1’s contextual misinterpretation, attributing the man’s smile to a general amusing memory rather than the dog’s barking resembling gunfire. Additionally, Gemini Pro 1.5 exhibits visually induced audio hallucination, mistaking the dog’s open-mouth appearance for barking and associating it with a crying baby.
@@
Figure D.6: Error analysis showing o1 and Gemini Pro 1.5 both struggling with spatial misjudgment and temporal misalignment, leading to errors in estimating the number of defenders due to difficulties in distinguishing actual close defenders and potential misalignment with the described action.
@@
Figure D.7: Error analysis showing Gemini Pro 1.5 providing inconsistent responses despite clear visuals and audio cues, failing to accurately identify the cause of the glass breaking due to rephrased question structure.
Appendix E Questionnaire Survey
@@
The survey included a total of 200 questions curated from the dataset, covering diverse subcategories within each defined split. For each participant, the questionnaire consisted of two parts: one multiple-choice question (MCQ) with eight answer options, followed by one open-ended question from a different video. To minimize content overlap and reduce potential bias, no video was reused between questions within a single questionnaire. This approach ensured broad coverage of the dataset’s taxonomy while avoiding contamination across tasks.
@@
We tested human performance across three different conditions: audio only, visual only, and audio plus video. Each condition was toggle-enabled within the survey toolkit, allowing participants to be assigned to a specific modality. The survey interfaces for the conditions are shown in Fig. E.2 for audio only, Fig. E.3 for visual only, Fig. E.1 for both modalities. To prevent cross-condition contamination and ensure focused evaluation, each participant completed the entire study under only one selected modality.
@@
Figure E.1: Screenshot of the questionnaire under the audio-plus-visual condition, showcasing the interface used to assess human performance across multiple-choice questions. Results offer a baseline for comparison with multimodal model performance across diverse tasks and difficulty levels.
Figure E.2: Questionnaire screenshot showing open-ended questions under the audio-only condition. Responses were evaluated using the GPT-4o grading pipeline, capturing the quality of language generation based solely on audio context.
Figure E.3: Questionnaire screenshot showing open-ended questions under the visual-only condition. GPT-4o evaluation scores assess human understanding based solely on visual information, without the aid of audio context.
Appendix F Dataset Distribution Continued
In this section, we provide a detailed breakdown of the taxonomy distributions within each defined agentic task category. For each task type, we analyze the composition of its associated taxonomies to illustrate the diversity and balance of the dataset across different functional scenarios. The corresponding distributions are visualized in Fig. F.1 for Gaming, Fig. F.2 for Egocentric Agent, Fig. F.3 for Information Querying, Fig. F.4 for Sentiments, Fig. F.5 for Shopping, Fig. F.6 for Social, and Fig. F.7 for Sports. These visualizations offer insight into the internal structure of each agentic task and highlight the variety of situations represented within the dataset.
Figure F.1: Taxonomy composition in Gaming.
Figure F.2: Taxonomy composition in Egocentric Agent.
Figure F.3: Taxonomy composition in Information Querying.
Figure F.4: Taxonomy composition in Sentiments.
Figure F.5: Taxonomy composition in Shopping.
Figure F.6: Taxonomy composition in Social.
Figure F.7: Taxonomy composition in Sports.

### v2

@@ C.6 Common Error Modes
To better understand the limitations of the state of the art models, we analyze common error patterns observed across the benchmark. We find that model failures often arise from modality-specific weaknesses, such as misinterpretation of complex visual cues, inability to process nuanced audio signals, or challenges in aligning multimodal information with nuanced context. For illustration, we highlight representative failure cases from OpenAI o1 and Gemini 1.5 Pro, chosen because they are among the strongest available models; however, the same error patterns recur across the models we evaluated. Common failure modes are showcased in Fig. 5,  6,  7,  8,  9,  10.
@@
Figure 5: Error analysis showing that o1 fails to correctly answer the question when the audio cannot be transcribed into text-based subtitles, leading to an incorrect connection between the cityscape video and the instrumental music.
@@
Figure 6: Error analysis showing inconsistent responses from o1, where it failed to accurately identify Hika-chan, when the question was rephrased slightly.
@@
Figure 7: Error analysis highlighting o1’s failure in correctly synchronizing the audio with visual cues, leading to misinterpretation of the woman picking up tree ornaments.
@@
Figure 8: Error analysis highlighting o1’s contextual misinterpretation, attributing the man’s smile to a general amusing memory rather than the dog’s barking resembling gunfire. Additionally, Gemini Pro 1.5 exhibits visually induced audio hallucination, mistaking the dog’s open-mouth appearance for barking and associating it with a crying baby.
@@
Figure 9: Error analysis showing o1 and Gemini Pro 1.5 both struggling with spatial misjudgment and temporal misalignment, leading to errors in estimating the number of defenders due to difficulties in distinguishing actual close defenders and potential misalignment with the described action.
@@
Figure 10: Error analysis showing Gemini Pro 1.5 providing inconsistent responses despite clear visuals and audio cues, failing to accurately identify the cause of the glass breaking due to rephrased question structure.
Appendix D D. Questionnaire Survey
@@
The survey included a total of 261 questions curated from the dataset, covering diverse subcategories within each defined split. For each participant, the questionnaire consisted of either one multiple-choice question (MCQ) with eight answer options, or an open-ended question from a different video. To minimize content overlap and reduce potential bias, no video was reused between questions within a single questionnaire. This approach ensured broad coverage of the dataset’s taxonomy while avoiding contamination across tasks. We recruited 382 participants through Amazon Mechanical Turk service with approval rate >>96% to answer the questions with A, S, V, A+V, S+V as inputs with media from MAVERIX and MAVERIX-Long. Both MCQ responses and open-ended responses are recorded.
@@
We tested human performance across three different conditions: audio only, visual only, and audio plus video. Each condition was toggle-enabled within the survey toolkit, allowing participants to be assigned to a specific modality. The survey interfaces for the conditions are shown in Fig. 12 for audio only, Fig. 13 for visual only, Fig. 11 for both modalities. To prevent cross-condition contamination and ensure focused evaluation, each participant completed the entire study under only one selected modality.
@@
Figure 11: Screenshot of the questionnaire under the audio-plus-visual condition, showcasing the interface used to assess human performance across multiple-choice questions. Results offer a baseline for comparison with multimodal model performance across diverse tasks and difficulty levels.
Figure 12: Questionnaire screenshot showing open-ended questions under the audio-only condition. Responses were evaluated using the GPT-4o grading pipeline, capturing the quality of language generation based solely on audio context.
Figure 13: Questionnaire screenshot showing open-ended questions under the visual-only condition. GPT-4o evaluation scores assess human understanding based solely on visual information, without the aid of audio context.
Appendix E E. Dataset Distribution Continued
In this section, we provide a detailed breakdown of the taxonomy distributions within each defined agentic task category. For each task type, we analyze the composition of its associated taxonomies to illustrate the diversity and balance of the dataset across different functional scenarios. The corresponding distributions are visualized in Fig. 14 for Gaming, Fig. 15 for Egocentric Agent, Fig. 16 for Information Querying, Fig. 17 for Sentiments, Fig. 18 for Shopping, Fig. 19 for Social, and Fig. 20 for Sports. These visualizations offer insight into the internal structure of each agentic task and highlight the variety of situations represented within the dataset.
Figure 14: Taxonomy composition in Gaming.
Figure 15: Taxonomy composition in Egocentric Agent.
Figure 16: Taxonomy composition in Information Querying.
Figure 17: Taxonomy composition in Sentiments.
Figure 18: Taxonomy composition in Shopping.
Figure 19: Taxonomy composition in Social.
Figure 20: Taxonomy composition in Sports.